OpenStudy (anonymous):

find the min of... g(t)=t^3-9t+5 on interval [-1.5, 3]

OpenStudy (anonymous):

why isnt it 3?

OpenStudy (anonymous):

don't i just take the derivative of whats given, set it to zero, and solve?

OpenStudy (anonymous):

g(t)=t^3-9t+5 on interval [-1.5, 3] g'(t)= 3t² - 9 => 3(3t - 3) => t = 1

OpenStudy (anonymous):

That's the first step, isn't it?

OpenStudy (anonymous):

oops i made a mistake

OpenStudy (anonymous):

3(t² - 3) = 0 => t = + sqrt3 and t = -sqrt 3

OpenStudy (anonymous):

did you do that?

OpenStudy (saifoo.khan):

yo!

OpenStudy (anonymous):

is that right, saifoo?

OpenStudy (anonymous):

i dont think it is

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (saifoo.khan):

Melind is correct

OpenStudy (anonymous):

i must not be putting it in my hw site correctly then...

OpenStudy (anonymous):

Now i'd probably do like this. Considering sqrt3 is 1,7 t -1,7 ... 1,5 ... 3 g'(t) g(t) and find the values and signs (positive and negative)

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

me too, kinda lol wait a sec pls

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

i guess we should draw a graph to see it clearer

OpenStudy (anonymous):

@KingGeorge hey you there?

OpenStudy (kinggeorge):

So you found the critical points at \(t= \pm \sqrt3\). Next step, take the second derivative. Then you get\[f''(t)=6t\]Since this is negative when \(t=-\sqrt3\), that must be where the minimum is. Last step, solve for \(f(-\sqrt3)\).

OpenStudy (kinggeorge):

That was a mistake. We're looking for where the second derivative is positive, not negative, so the minimum would be at \(t=+\sqrt3\). So solve for that instead of \(-\sqrt3\).

OpenStudy (kinggeorge):

You should get \[f(\sqrt3)=3\sqrt3 -9\sqrt3+5=-6\sqrt3+5\]

OpenStudy (anonymous):

you forgot the squared

OpenStudy (kinggeorge):

Now we need to check the boundaries. If we try \(t=-1.5\), we get a value greater than \(f(\sqrt3)\), and if we try \(t=3\), we also get a value greater than \(f(\sqrt3)\). So \(-6\sqrt3+5\) is the minimum value.

OpenStudy (kinggeorge):

The function is \[t^3-9t+5\] correct? In my calculation above I've already cubed \(t\)

OpenStudy (anonymous):

nvm then

OpenStudy (anonymous):

Thanks :D I posted another Max/Min prob if you wanna help with that too.

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (anonymous):

King is awesome, thanks.

OpenStudy (kinggeorge):

Why thank you. :)